# On the relation between cosmological redshift and fine structure constant variation

### v3

@@ 1 Introduction
Different cosmological tests were proposed to probe whether the Universe is expanding or remains static. Tolman (1930) predicted that in an expanding universe, the surface brightness of a receding source with redshift $z$ will be dimmed by $\sim(1+z)^{-4}$. Consequently to Tolman’s prediction, the equation $d_{L}=d_{A}(1+z)^{\gamma}$, with $\gamma=2$ was established between luminosity distance $d_{L}$ and angular diameter distance $d_{A}$ for a expanding universe. There are contradictory results on this kind of test, some of them claim for expansion (Lubin & Sandage (2001); Sandage (2010)) while others (Lerner (2006); Crawford (2010); Brynjolfsson (2006)) are advocating for a static universe. Nevertheless, neither expanding nor static studies are conclusive, since they depend on an uncertain possible galaxy evolution. In this work, we develop the Cosmic Time Physics theory that offer an alternative explanation to cosmological redshift, conducting to a new consistent cosmological model.
@@ 2 Foundations of the expanding universe
Eq.  19 is commonly known as Etherington distance-duality relation.

### v4

# Cosmic Time Physics – On the relation between cosmological redshift and fine structure constant variation

@@ 1 Introduction
Different cosmological tests were proposed to probe whether the Universe is expanding or remains static. Tolman (1930) predicted that in an expanding universe, the surface brightness of a receding source with redshift $z$ will be dimmed by $\sim(1+z)^{-4}$. Consequently to Tolman’s prediction, the equation $d_{L}=d_{A}(1+z)^{\gamma}$, with $\gamma=2$ was established between luminosity distance $d_{L}$ and angular diameter distance $d_{A}$ for a expanding universe. There are contradictory results on this kind of test, some of them claim for expansion (Lubin and Sandage (2001); Sandage (2010)) while others (Lerner (2006); Crawford (2010); Brynjolfsson (2006)) are advocating for a static universe. Nevertheless, neither expanding nor static studies are conclusive, since they depend on an uncertain possible galaxy evolution. In this work, we develop the Cosmic Time Physics theory that offer an alternative explanation to cosmological redshift, conducting to a new consistent cosmological model.
@@ 2 Foundations of the expanding universe
Eq.  19 is commonly known as Etherington distance-duality relation. It is important to clarify that the original Etherington (1933) equation is the reciprocity theorem for a local (i.e. non-expanding) universe:
$d_{L}=d_{A}(1+z)\qquad(non-expanding\ universe)$ (20)
It was reformulated for an expanding universe by adding another $(1+z)$ factor.
@@ 4 Discussion
Finally, let us to reveal the illusion of the expansing universe corresponding to the FLRW metric. Dividing the FLRW metric (Eq. 2) by $a(t)^{2}$ we have
$-c^{2}\frac{d\tau^{2}}{a(t)^{2}}=-c^{2}\frac{dt^{2}}{a(t)^{2}}+\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}d\Omega^{2}\right]$ (64)
Now, making the change of variables, $dt^{\prime}=dt/a(t)$ and $d\tau^{\prime}=d\tau/a(t)$, the Eq. 64 can be expressed as
$-c^{2}d\tau^{\prime 2}=-c^{2}dt^{\prime 2}+\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}d\Omega^{2}\right]$ (65)
where the scale factor $a(t)$ disappear, fading the expanding universe illusion. Note, that the FLRW is still valid for a homogeneous and isotropic non-expanding universe.